# ArtNet: Hierarchical Clustering-Based Artificial Netlist Generator for ML and DTCO Applications

### v1

@@ IV-B1 Convergence
We highlight both maximum (blue) and minimum (red) values for the target parameter, while tracking the evolution of all parameters throughout the convergence process from start (dashed lines) to end (solid lines). The three main columns in Fig. 7 show how extreme values in one parameter (e.g., maximum $p=0.80$ in blue, minimum $p=0.30$ in pink, and purple overlap in the left column) affect values taken by other parameters during the convergence process. The narrower bands for ArtNet show its superior convergence behavior as well as smaller impact of any given parameter on other parameters, relative to GNL and ANG. We see that ArtNet not only converges more reliably toward the requested parameter values, but also minimizes cross-parameter interference, which is critical for predictable netlist generation.
Fig. 7: Convergence plot. The $x$, $y$, $z$, $w$ axes represent the number of instances, $p$, $S_{ratio}$, and $N_{PI}+N_{PO}$, with ranges of [0, 20,000], [0.30, 0.80], [0, 0.40], [0, 1000], respectively.
@@ IV-B2 Coverage
Data
Points
Excluded $S_{ratio}$ $p$ #Insts
@@
Fig. 8: Coverage plot. The number of instances, $S_{ratio}$, and $p$ have ranges of [10,000, 200,000], [0.05, 0.30], and [0.35, 0.70], respectively.
We evaluate the ability of ANG, GNL, and ArtNet to achieve broad parameter space coverage thorough exploration of diverse design configurations. Key parameters include #Insts (design space), $S_{ratio}$ (timing path), and $p$ (interconnect complexity). For ArtNet and GNL, we sweep #Insts within [10,000, 200,000] with step size 10,000; we sweep $S_{ratio}$ within [0.05, 0.30] with step size 0.05; and we sweep $p$ within [0.35, 0.70] with step size 0.05. Since ANG does not use $p$ as an input, we sweep alternative parameters related to routing complexity: average net degree within [2.0, 3.0] with step size 0.5, and average bounding box size within [0.1, 1.0] with step size 0.3. This setup results in 1120 data points each for ArtNet and GNL, and 1280 data points for ANG. Table IV shows the netlist parameter coverage data and Fig. 8 visualizes the coverages across the parameter space, where parameter combinations that induce artificial netlists outside of the space are considered to be outliers and excluded from the figure. The excluded outliers comprise 766 points from ANG (mostly due to high $p$ values), 421 points from GNL (due to high $S_{ratio}$), and 145 points from ArtNet (due to $S_{ratio}$ deviations and low $p$). As the design size increases, ANG shows increasing $p$, and struggles to maintain uniform $S_{ratio}$. For GNL, additional FFs frequently push $S_{ratio}$ beyond the limit, and also cause instance count limits to be exceeded. In contrast, ArtNet shows more stable performance across the range of input parameter combinations, and demonstrates superior coverage.
@@ IV-C2 Rent’s Analysis
*The value closest to Real (target) is highlighted in Bold. ArtNet “wins” $36\frac{1}{3}$ of 72 (4 methods $\times$ 3 types $\times$ 6 metrics) head-to-head
comparisons against ANG and GNL, where $k$-way ties are counted as $\frac{1}{k}$ fractional wins.
We evaluate Rent’s parameter using these 24 different methods, combining four evaluation methods, three pin counting methods, and two averaging methods. This comprehensive study provides a thorough analysis across structural and placement patterns, giving a clear picture of how artificial netlists compare to real designs. We evaluate Rent’s exponent on six designs listed in Table VIII. Table VII summarizes the errors between Rent’s exponent values of real and artificial designs using mean absolute percentage error (MAPE), mean absolute error (MAE), and median absolute error (MedAE). ArtNet demonstrates consistently low error in estimating Rent’s parameters across all evaluation methods, with particularly strong performance for type2 pins. In circuit partitioning method, ArtNet achieves a MAPE of 1.47%, MAE of 0.01, and MedAE of 0.01, significantly outperforming ANG and GNL. In graph traversal, ArtNet maintains superior accuracy with a MAPE of 3.43%, MAE of 0.026, and MedAE of 0.02. Similarly, in rectangle sampling I, ArtNet achieves a MAPE of 5.24%, MAE of 0.026, and MedAE of 0.02, while in rectangle sampling II, it achieves a MAPE of 3.64%, MAE of 0.02, and MedAE of 0.01, demonstrating consistent and precise estimation across evaluation methods. In particular, ArtNet shows significantly lower errors in the circuit partitioning-based method because its recursive clustering closely aligns with the partitioning approach. We see that the netlists generated by ArtNet closely match not only the structural characteristics but also the placement patterns of the real designs.
@@
  Pre-Place Post-Place Post-Route
Runtime
Design #Insts #Nets #$F$s $p$ HPWL PinD rWL WNS TNS Power #DRVs
jpeg Real 46,593 46,612 4,392 0.49 1.000 0.314 1.000 -37 -5.61 0.071 0 -
ANG 45,627 45,645 8,741 0.73 4.089 0.236 4.176 -114 -157.2 0.109 1,028 257.3
GNL 50,569 50,571 8,368 0.54 1.327 0.272 1.366 -16 -1.175 0.104 0 0.41
ArtNet 46,682 46,703 4,482 0.50 1.054 0.309 1.013 -39 -0.816 0.069 0 0.23
nova Real 142,863 143,261 29,131 0.57 1.000 0.306 1.000 -122 -24.64 0.059 6 -
ANG 120,704 121,169 25,890 0.76 2.062 0.209 2.121 -22 -1.087 0.092 24 3,160
GNL 145,720 146,409 31,988 0.59 1.297 0.293 1.347 -1466 -27.7k 0.155 0 1.82
ArtNet 142,547 143,236 28,815 0.56 1.088 0.299 1.015 -367 -42.03 0.071 0 0.76
netcard Real 261,591 263,416 67,440 0.58 1.000 0.251 1.000 -28 -0.574 0.341 5 -
ANG 238,861 240,700 59,926 0.81 4.290 0.209 N/A N/A N/A N/A N/A 18,427
GNL 270,180 272,024 76,029 0.60 1.133 0.227 1.037 0.00 0.00 0.575 0 3.95
ArtNet 260,113 261,957 65,962 0.57 1.113 0.242 0.983 -4 -0.01 0.359 0 1.61
tate_pairing Real 211,468 212,246 31,416 0.54 1.000 0.348 1.000 -20 -1.451 0.303 0 -
ANG 201,401 202,179 44,716 0.75 5.987 0.265 N/A N/A N/A N/A N/A 3525
GNL 223,663 224,442 43,611 0.56 1.763 0.305 1.779 -161 -1419.5 0.543 1 3.14
ArtNet 211,177 211,956 31,125 0.54 1.453 0.338 1.406 -11 -0.265 0.322 1 1.47
CA53 Real 303,242 305,294 35,234 0.65 1.000 0.098 1.000 -104 -94.26 0.392 202 -
ANG N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A -
GNL 317,373 320,006 49,365 0.66 1.785 0.092 1.727 -610 -13k 0.827 9,734 20.34
ArtNet 303,251 306,235 35,243 0.63 0.985 0.096 0.961 -290 -317.3 0.431 649 13.49
ariane-133 Real 97,329 100,000 19,807 0.58 1.000 0.056 1.000 -21 -2.54 0.341 3 -
ANG N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A -
GNL 97,755 101,625 19,807 0.59 1.376 0.058 1.385 -7 -0.015 0.405 14 2.06
ArtNet 97,496 101,051 19,975 0.58 1.111 0.058 1.091 0.00 0.00 0.342 0 0.69
*WNS is in units of $ps$; TNS is in units of $ns$; and Power is in units of $W$. The value closest to Real (target) is highlighted in Bold.
@@ IV-C3 P&R Evaluation
Topological features such as circuit size, interconnect complexity, and timing depth significantly influence P&R behavior. We experimentally confirm that ArtNet’s artificial netlists designed to mimic real circuit topologies can achieve comparable P&R QoR metrics, validating their use as reliable proxies for design benchmarking. For our experiments, we use jpeg, nova, netcard, tate_pairing, Arm Cortex-A53 (CA53), and ariane-133 [39]. Table VIII shows the QoR metrics for each circuit from real, ANG, GNL, and ArtNet versions. All evaluations use identical design parameters and the same tool flow for consistency.
We compare artificial netlists generated by ANG, GNL and ArtNet across three design stages. At pre-placement, we evaluate the number of instances (#Insts), number of nets (#Nets), number of flip-flops (#FFs) and Rent’s exponent ($p$). At post-placement, we evaluate half-perimeter wire length (HPWL) and pin density [^4] Pin density is defined as the number of pins divided by the core area. (PinD). At the post-routing stage, we evaluate routed wirelength (rWL), worst negative slack (WNS), total negative slack (TNS), total power (Power), and the number of DRVs (#DRVs). We note that ANG fails to generate designs for CA53 and ariane-133 due to its lack of macro insertion support. Also, while ANG can process netcard and tate_pairing parameters, the ANG-produced netlists fail routing due to excessive interconnect complexity (>> 8.5M DRVs and >> 3M DRVs, respectively).
@@
Table X compares the changes in area and #Insts during the downstream optimization design process, assuming that a well-replicated netlist should exhibit similar design feature changes to those in real designs. Each delta value is calculated by subtracting the pre-place stage value from the post-route optimization stage value and dividing the result by the pre-place stage value. On average across six designs, ArtNet undergoes a 4.30% change in area and a 5.90% change in #Insts, showing the closest alignment with the changes observed in the real design.
@@ IV-C4 Design Locality
Fig. 9 illustrates ArtNet’s ability to reflect both structural and timing-related locality of real design. We use ArtNet to generate two netlists: (1) ArtNet-generated nova with logical hierarchy, and (2) ArtNet-generated nova without logical hierarchy. Then, we compare these netlists on four different aspects: clustering results; timing critical paths; routing congestion map; and static IR drop.
@@
Next, we analyze the routing congestion map. Fig. 9(c) shows the congestion distribution across metal layers M1 to M4. In both the real and hierarchy-aware netlists, congestion is concentrated in specific regions. In contrast, the flat netlist shows a more uniform congestion spread, indicating a loss of spatial locality. Last, we examine the IR drop behavior after routing. Fig. 9(d) shows the IR drop distribution. In both the real and hierarchy-aware netlists, IR drop hotspots can be observed. The flat netlist, however, shows more diffuse IR drop and a loss of locality.
@@ IV-D Assessment in an ML Framework
We demonstrate a practical use case of artificial netlists in the ML context. We generate augmented datasets using ANG and ArtNet, apply these datasets in the training of ML models, evaluate model performance on unseen circuits, and compare against a baseline trained on a real dataset.
Design of Experiments. We train three CNN-based routability prediction models [25, 38] respectively using three datasets: (1) $\text{M}_{\text{real}}$ with real netlist dataset, (2) $\text{M}_{\text{ANG}}$ with augmented dataset from ANG, and (3) $\text{M}_{\text{ArtNet}}$ with augmented dataset from ArtNet. For the real design dataset, we generate 48 layouts per design, varying utilization, top routing layer, and clock period. The real designs [34] used are as follows:
  • Training dataset. aes_cipher_top, eth_top, ibex_core, mpeg2_top, point_scalar_mult, tate_pairing, vga_enh_top, des3, fpu, keccak, pci_bridge32, sasc_top, tv80s, wb_conmax_top
  • Test dataset. ldpc_decoder_802_3an, netcard, nova
To create the augmented datasets, we use the Sobol sequence [27] for efficient parameter sampling with broad coverage of the parameter space. Table XIII provides the parameter space, determined based on the real design training dataset. We sample 42 configurations and generate 42 netlists, each with 16 layouts that vary in utilization, top routing layer, and clock period, for a total of 672 layouts. The final augmented dataset combines both real and artificial datasets.
@@
Model Performance Evaluation. Table XIV compares the models using accuracy, recall, precision, and F1 score. The $\text{M}_{\text{real}}$ model achieves 0.93 accuracy, 0.46 recall, 0.67 precision, and F1 score of 0.56. With the benefit of data augmentation, the $\text{M}_{\text{ANG}}$ model improves all four metrics. The $\text{M}_{\text{ArtNet}}$ model achieves the best results, with 0.96 accuracy, 0.73 recall, 0.72 precision, and F1 score of 0.72.
@@
Mreal 0.93 0.46 0.67 0.56
MANG 0.94 0.61 0.71 0.66
MArtNet 0.96 0.73 0.72 0.72
Dataset Analysis. Table XV presents a statistical comparison between the ANG and ArtNet datasets based on four key metrics: variance [11], mean pairwise distance [3], and mean and standard deviation of eigenvalues of the CNN feature map (treated as a matrix, following [13]). We calculate these metrics from data points within a single layout to reflect the feature diversity of each layout. Larger variance and higher mean pairwise distance indicate greater sample diversity. Higher mean eigenvalues suggest more variability captured by principal components, while a lower standard deviation implies a more even distribution.
@@
Fig. 11 illustrates feature diversity for both datasets using a principal component analysis (PCA) plot [13]. The ANG dataset exhibits a clustered and skewed distribution, indicating limited variability and potential feature redundancy. In contrast, ArtNet dataset presents a more uniformly dispersed distribution that spans a wider area of the feature space. Such a broader distribution implies richer structural information, which is consistent with extraction of more effective features that ultimately result in better model performance [10].

### v2

@@ IV-B1 Convergence
We highlight both maximum (blue) and minimum (red) values for the target parameter, while tracking the evolution of all parameters throughout the convergence process from start (dashed lines) to end (solid lines). The three main columns in Fig. 7 show how extreme values in one parameter (e.g., maximum $p=0.80$ in blue, minimum $p=0.30$ in red, and purple overlap in the left column) affect values taken by other parameters during the convergence process. The narrower bands for ArtNet show its superior convergence behavior as well as smaller impact of any given parameter on other parameters, relative to GNL and ANG. We see that ArtNet not only converges more reliably toward the requested parameter values, but also minimizes cross-parameter interference, which is critical for predictable netlist generation.
Fig. 7: Convergence plot. The ranges of $N_{inst}$, $p$, $S_{ratio}$, and $N_{PIO}$ (= $N_{PI}$ + $N_{PO}$ ) are [0, 20,000], [0.30, 0.80], [0, 0.40], and [0, 1000], respectively (Refer to Table III for notation.)
@@ IV-B2 Coverage
  Data Points Excluded $S_{ratio}$ $p$ #Insts
@@
Fig. 8: Coverage plot. The number of instances, $S_{ratio}$, and $p$ have ranges of [10,000, 200,000], [0.05, 0.30], and [0.35, 0.70], respectively. (Refer to Table III for notation.)
We evaluate the ability of ANG, GNL, and ArtNet to achieve broad parameter space coverage through exploration of diverse design configurations. Key parameters include #Insts (design space), $S_{ratio}$ (timing path), and $p$ (interconnect complexity). For ArtNet and GNL, we sweep #Insts within [10,000, 200,000] with step size 10,000; we sweep $S_{ratio}$ within [0.05, 0.30] with step size 0.05; and we sweep $p$ within [0.35, 0.70] with step size 0.05. Since ANG does not use $p$ as an input, we sweep alternative parameters related to routing complexity: average net degree within [2.0, 3.0] with step size 0.5, and average bounding box size within [0.1, 1.0] with step size 0.3. This setup results in 1120 data points each for ArtNet and GNL, and 1280 data points for ANG. Table IV shows the netlist parameter coverage data and Fig. 8 visualizes the coverages across the parameter space, where parameter combinations that induce artificial netlists outside of the space are considered to be outliers and excluded from the figure. The excluded outliers comprise 766 points from ANG (mostly due to high $p$ values), 421 points from GNL (due to high $S_{ratio}$), and 145 points from ArtNet (due to $S_{ratio}$ deviations and low $p$). As the design size increases, ANG shows increasing $p$, and struggles to maintain uniform $S_{ratio}$. For GNL, additional FFs frequently push $S_{ratio}$ beyond the limit, and also cause instance count limits to be exceeded. In contrast, ArtNet shows more stable performance across the range of input parameter combinations, and demonstrates superior coverage. This broad coverage reflects diversity with reduced bias obtained by sweeping input parameters, while accuracy is preserved during construction.
@@ IV-C2 Rent’s Analysis
We evaluate Rent’s parameter using these 24 different methods, combining four evaluation methods, three pin counting methods, and two averaging methods. This comprehensive study provides a thorough analysis across structural and placement patterns, giving a clear picture of how artificial netlists compare to real designs. We evaluate Rent’s exponent on six designs listed in Table VIII. Table VII summarizes the errors between Rent’s exponent values of real and artificial designs using mean absolute percentage error (MAPE), mean absolute error (MAE), and median absolute error (MedAE). ArtNet demonstrates consistently low error across all evaluation methods , with particularly strong performance for Type2 pin counting, as it adopts a hyperedge-based formulation that directly models net-level connectivity. Type1 and Type3 results are reported as complementary graph-edge views. In circuit partitioning method, ArtNet achieves a MAPE of 1.43%, MAE of 0.01, and MedAE of 0.01, significantly outperforming ANG and GNL. In graph traversal, ArtNet maintains superior accuracy with a MAPE of 3.43%, MAE of 0.026, and MedAE of 0.02. Similarly, in rectangle sampling I, ArtNet achieves a MAPE of 5.24%, MAE of 0.026, and MedAE of 0.02, while in rectangle sampling II, it achieves a MAPE of 3.64%, MAE of 0.02, and MedAE of 0.01, demonstrating consistent and precise estimation across evaluation methods. In particular, ArtNet shows significantly lower errors in the circuit partitioning-based method because its recursive clustering closely aligns with the partitioning approach. We see that the netlists generated by ArtNet closely match not only the structural characteristics but also the placement patterns of the real designs.
Differences across pin types are largely attributable to the generators’ construction assumptions. ANG builds nets from graph-edge–level connections, which tends to favor Type1 metrics, whereas ArtNet and GNL generate nets directly at the hyperedge level, aligning more closely with Type2 evaluation. Consistent with this, ArtNet attains the lowest error in all 24 Type2 settings and remains competitive under graph-edge metrics. Overall, these results indicate that the netlists generated by ArtNet closely match the net-level interconnect structure of real netlists, while also showing robust behavior under alternative graph-edge–based evaluations.
@@
*The value closest to Real (target) is highlighted in Bold. ArtNet “wins” $36\frac{1}{3}$ of 72 (4 methods $\times$ 3 types $\times$ 6 metrics) head-to-head comparisons against ANG ( $21\frac{5}{6}$) and GNL ( $13\frac{5}{6}$), where $k$-way ties are counted as $\frac{1}{k}$ fractional wins.
@@
  Pre-Place Post-Place Post-Route Runtime (s)
Design #Insts #Nets #$F$s $p$ HPWL PinD Area CTS Leak rWL WNS TNS Power #DRVs
jpeg Real 46,593 46,612 4,392 0.49 1.000 0.314 5,456 7.87 1.000 -37 -5.61 0.071 0 -
ANG 45,627 45,645 8,741 0.73 4.089 0.236 6,353 12.68 4.176 -114 -157.2 0.109 1,028 257.3
GNL 50,569 50,571 8,368 0.54 1.327 0.272 7,307 9.96 1.366 -16 -1.175 0.104 0 0.41
ArtNet 46,682 46,703 4,482 0.50 1.054 0.309 5,620 5.29 1.013 -39 -0.816 0.069 0 0.23
nova Real 142,863 143,261 29,131 0.57 1.000 0.306 19,037 31.81 1.000 -122 -24.64 0.059 6 -
ANG 120,704 121,169 25,890 0.76 2.062 0.209 15,019 39.40 2.121 -22 -1.087 0.092 24 3,160
GNL 145,720 146,409 31,988 0.59 1.297 0.293 24,204 37.32 1.347 -1466 -27.7k 0.155 0 1.82
ArtNet 142,547 143,236 28,815 0.56 1.088 0.299 17,257 32.58 1.015 -367 -42.03 0.071 0 0.76
netcard Real 261,591 263,416 67,440 0.58 1.000 0.251 42,488 75.04 1.000 -28 -0.574 0.341 5 -
ANG 238,861 240,700 59,926 0.81 4.290 0.209 N/A N/A N/A N/A N/A N/A >> 8.5M 18,427
GNL 270,180 272,024 76,029 0.60 1.133 0.227 47,447 80.81 1.037 0.00 0.00 0.575 0 3.95
ArtNet 260,113 261,957 65,962 0.57 1.113 0.242 38,073 71.34 0.983 -4 -0.01 0.359 0 1.61
tate_pairing Real 211,468 212,246 31,416 0.54 1.000 0.348 23,772 36.97 1.000 -20 -1.451 0.303 0 -
ANG 201,401 202,179 44,716 0.75 5.987 0.265 N/A N/A N/A N/A N/A N/A >> 3M 3525
GNL 223,663 224,442 43,611 0.56 1.763 0.305 34,149 49.19 1.779 -161 -1419.5 0.543 1 3.14
ArtNet 211,177 211,956 31,125 0.54 1.453 0.338 24,655 35.77 1.406 -11 -0.265 0.322 1 1.47
CA53 Real 303,242 305,294 35,234 0.65 1.000 0.098 39,746 48.37 1.000 -104 -94.26 0.392 202 -
ANG N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A -
GNL 317,373 320,006 49,365 0.66 1.785 0.092 63,630 73.02 1.727 -610 -13k 0.827 9,734 20.34
ArtNet 303,251 306,235 35,243 0.63 0.985 0.096 35,974 51.85 0.961 -290 -317.3 0.431 649 13.49
ariane-133 Real 97,329 100,000 19,807 0.58 1.000 0.056 15,909 0.38 1.000 -21 -2.54 0.341 3 -
ANG N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A -
GNL 97,755 101,625 19,807 0.59 1.376 0.058 19,457 0.35 1.385 -7 -0.015 0.405 14 2.06
ArtNet 97,496 101,051 19,975 0.58 1.111 0.058 14,618 0.35 1.091 0.00 0.00 0.342 0 0.69
*Area is in units of $\mu m^{2}$; CTS Leak is in units of $\mu W$; WNS is in units of $ps$; TNS is in units of $ns$; and Power is in units of $W$.
The value closest to Real (target) is highlighted in Bold.
@@ IV-C3 P&R Evaluation
Topological features such as circuit size, interconnect complexity, and timing depth significantly influence P&R behavior. We experimentally confirm that ArtNet’s artificial netlists designed to mimic real circuit topologies can achieve comparable P&R QoR metrics, validating their use as reliable proxies for design benchmarking. For our experiments, we use jpeg, nova, netcard, tate_pairing, Arm Cortex-A53 (CA53), and ariane-133 [41]. Table VIII shows the QoR metrics for each circuit from real, ANG, GNL, and ArtNet versions after running commercial P&R flow. All evaluations use identical design parameters and the same tool flow for consistency.
We compare artificial netlists generated by ANG, GNL and ArtNet across three design stages. At pre-placement, we evaluate the number of instances (#Insts), number of nets (#Nets), number of flip-flops (#FFs) and Rent’s exponent ($p$) , directly extracted from the netlists. At post-placement, we evaluate half-perimeter wire length (HPWL) and pin density [^4] Pin density is defined as the number of pins divided by the core area. (PinD). At the post-routing stage, we evaluate routed wirelength (rWL), worst negative slack (WNS), total negative slack (TNS), total power (Power), and the number of DRVs (#DRVs). We note that ANG fails to generate designs for CA53 and ariane-133 due to its lack of macro insertion support. Also, while ANG can process netcard and tate_pairing parameters, the ANG-produced netlists fail routing due to excessive interconnect complexity (>> 8.5M DRVs and >> 3M DRVs, respectively).
@@
Table X compares the changes in area and #Insts during the downstream optimization design process, assuming that a well-replicated netlist should exhibit similar design feature changes to those in real designs , as well as similar design feature values in Table VIII. Each delta value is calculated by subtracting the pre-place stage value from the post-route optimization stage value and dividing the result by the pre-place stage value. On average across six designs, ArtNet undergoes a 4.30% change in area and a 5.90% change in #Insts, showing the closest alignment with the changes observed in the real design. Although ANG achieves better results than ArtNet for jpeg, Table VIII shows that ArtNet, overall, generates netlists more similar to the real design.
@@ IV-C4 Design Locality
Fig. 9 illustrates ArtNet’s ability to reflect both structural and timing-related locality of real design. While logical hierarchy-aware generation improves locality over flat generation, differences from real designs remain. We use ArtNet to generate two netlists: (1) ArtNet-generated nova with logical hierarchy, and (2) ArtNet-generated nova without logical hierarchy. Then, we compare these netlists on four different aspects: clustering results; timing critical paths; routing congestion map; and static IR drop.
@@
Next, we analyze the routing congestion map. Fig. 9(c) shows the congestion distribution across metal layers M1 to M4. In both the real and hierarchy-aware netlists, congestion is concentrated in specific regions. In contrast, the flat netlist shows a more uniform congestion spread, indicating a loss of spatial locality. Last, we examine the IR drop behavior after routing. Fig. 9(d) shows the IR drop distribution. In both the real and hierarchy-aware netlists, IR drop hotspots can be observed. The flat netlist, however, shows more diffuse IR drop and a loss of locality. While ArtNet captures several key aspects of design locality, we acknowledge that noticeable differences remain compared to real designs.
@@ IV-D Assessment in an ML Framework
We demonstrate a practical use case of artificial netlists in the ML context. Fig. 10 illustrates the data augmentation flow with netlist generator. We generate augmented datasets using ANG and ArtNet, apply these datasets in the training of ML models, evaluate model performance on unseen circuits, and compare against a baseline trained on a real dataset.
Design of Experiments. We train three CNN-based routability prediction models [27, 40] respectively using three datasets: [^5] We adopt the same ML model used in ANG [22] for a fair comparison. (1) $\text{M}_{\text{real}}$ with real netlist dataset, (2) $\text{M}_{\text{ANG}}$ with augmented dataset from ANG, and (3) $\text{M}_{\text{ArtNet}}$ with augmented dataset from ArtNet. For the real design dataset, we generate 64 layouts per design, varying utilization, top routing layer, clock period , and aspect ratio. The real designs [36] used are as follows:
  • Training dataset. ldpc_decoder_802_3an, eth_top, ibex_core, mpeg2_top, point_scalar_mult, tate_pairing, vga_enh_top, des3, fpu, keccak, pci_bridge32, sasc_top, tv80s, wb_conmax_top
  • Test dataset. aes_cipher_top, netcard, nova
To create the augmented datasets, we use the Sobol sequence [29] for efficient parameter sampling with broad coverage of the parameter space. Table XIII provides the parameter space, determined based on the real design training dataset. We sample 64 configurations and generate 64 netlists, each with 18 layouts that vary in utilization, top routing layer, clock period , and aspect ratio, for a total of 1,152 layouts. The final augmented dataset combines both real and artificial datasets. Due to the tile-and-crop training pipeline of the CNN-based model, each layout is decomposed into many tiles, resulting in an effective training set that is much larger than the number of layouts and sufficient for mini-batch training.
@@
Model Performance Evaluation. Table XIV compares the models using accuracy, recall, precision, and F1 score. The $\text{M}_{\text{real}}$ model achieves 0.93 accuracy, 0.55 recall, 0.67 precision, and F1 score of 0.61. With the benefit of data augmentation, the $\text{M}_{\text{ANG}}$ model improves all four metrics. The $\text{M}_{\text{ArtNet}}$ model achieves the best results, with 0.96 accuracy, 0.72 recall, 0.76 precision, and F1 score of 0.74. For completeness, we further conduct a data-ablation study by training with 25%, 50%, 75%, and 100% of the ArtNet-augmented dataset, and with an ArtNet-only setting in which training uses only ArtNet-generated layouts (i.e., excluding all real-layout samples). We observe that performance largely saturates once 50% of the ArtNet-augmented data is used, and the ArtNet-only model remains comparable to the real-data baseline.
@@
Mreal 0.93 0.55 0.67 0.61
MANG 0.94 0.68 0.74 0.71
MArtNet 0.96 0.72 0.76 0.74
MArtNet_75% 0.96 0.71 0.77 0.73
MArtNet_50% 0.94 0.75 0.70 0.71
MArtNet_25% 0.93 0.60 0.69 0.64
MArtNet_only 0.92 0.58 0.60 0.59
Dataset Analysis. Table XV presents a statistical comparison between the ANG and ArtNet datasets based on four key metrics: variance ($\sigma^{2}$) [13], mean pairwise distance ($\bar{d}$) [3], and mean ($\bar{\lambda}$) and standard deviation of eigenvalues ($\text{STD}(\lambda)$) of the CNN feature map (treated as a matrix, following [15]). We calculate these metrics from data points within a single layout to reflect the feature diversity. Larger variance and higher mean pairwise distance indicate greater sample diversity. Higher mean eigenvalues suggest more variability captured by principal components, as the sum of the eigenvalues equals the total variance in the dataset. A lower standard deviation implies a more even distribution , as eigenvalues of roughly the same size mean the variance is spread out evenly.
@@
Fig. 11 illustrates feature diversity for both datasets using a principal component analysis (PCA) plot [15]. The ANG dataset exhibits a clustered and skewed distribution, indicating limited variability and potential feature redundancy. In contrast, ArtNet dataset presents a more uniformly dispersed distribution that spans a wider area of the feature space, although some skewness remains due to the inherent correlations among layout metrics. Such a broader distribution implies richer structural information, which is consistent with extraction of more effective features that ultimately result in better model performance [12].